A PARTICLE MOVES IN A STRAIGHT LINE AND ITS POSITION X AT TIME T IS GIVEN BY X2 =2+T ITS ACC. GIVEN

Grade:12th pass

1 Answers

Antara
2 years ago
x^2=2+T
Differentiating both sides,
2xdx=dT
So, v= dx/dT=1/2x
acceleration(a)=dv/dT=(dv/dx)*(dx/dT) ….(Chain rule)